# Graph RAG for Legal Norms: A Hierarchical and Temporal Approach

### v1

@@ 3.2 Versions as Aggregations
In Figure 11, we observe that on the norm’s original date (1988-10-05), the original version of the component "tit2" is formed by the original versions of its children. That is, the text of "Título II” (Title II) on that date consists of the texts of its children from that same date.
On 1993-09-14, a new version for “tit2 cap4" was produced, either by its own alteration or indirectly by the alteration of one or more components in its tree. A more common implementation would be to produce new versions for other children of "tit2" on 1993-09-14 and then produce a 1993-09-14 version for "tit2", which would be a composition [^16] In UML terms, Composition represents a strong whole-part relationship where the ”part” belongs exclusively to one ”whole” and shares its lifecycle; if the whole is deleted, the part is too. In our initial comparison, this would mean a parent version strictly owns child versions from the exact same date. of its children’s versions on the same date - similar to the formation of the "tit2" text on the original date.
@@
We can verify the formation of these version aggregations in Figure 11, where new versions of the component "tit2" were produced as a consequence of new versions of one or more of its children on each date.
@@
When a version of the child component Cx does not exist for a date Tn, the most recent version of Cx, prior to Tn, will be used in the aggregation, as exemplified by V1 of C3, which is being used in the formation of version V2 of C1 (examples of this situation also appear in Figure 11). In this case, Version V1 of component C3 contributes to version V1 of component C1 and to version V2 of component C1, demonstrating that a child version can participate in multiple parent versions in different temporal contexts. The orange arrows indicate the aggregation relationship (Figure 12).
@@
In the context of RAG, it allows the system to retrieve not only the specific version of a component, but also the context of its parent and child versions, providing a more holistic and precise understanding of the norm at a given moment, facilitating more complex queries about the temporal evolution of norms [2] - for example, a legal researcher will be able to ask what the wording of a certain article was on a specific date and obtain the precise answer thanks to the temporal component of Graph RAG.
@@ 3.3 Additional Context for Versions
For example, consider the caput of Art. 6º (presented in Figure 5) of the “Constituição Federal Brasileira de 1988” (Brazilian Federal Constitution of 1988), signed and published on 1988-10-05. Thus, at the time of the Constitution’s publication, the Additional Context property of its original version (1988-10-05) will contain information about the start of its validity:
@@
For example, with the publication, on 2000-02-15, of “Emenda Constitucional nº 26 de 14/02/2000” (Brazilian Constitutional Amendment No. 26), signed on 2000-02-14, an action was commanded, through the caput of its Art. 1º, to alter the wording of the caput provision of Art. 6° of the Brazilian Federal Constitution of 1988 – the action is represented in the graph [^19] The modeling of the “Action presented in Figure 13 was simplified for didactic purposes. It can be improved using ontology techniques, such as those proposed in UFO (Unified Foundational Ontology), found at https://nemo.inf.ufes.br/projetos/ufo/. Accessed on 22.03.2025. by the Action [^20] For a more precise representation in the Graph, specialized entity types can be created, derived from Action, for each type of action between norms found in the treated legal system. For example, in the Brazilian case, types like ”Alteração de Texto” (Text Change), ”Renumeração” (Renumbering), ”Revogação” (Repeal) etc. Thus, the entity representing the action could be included in the graph using the specialized type. entity (in brown in Figure 13). This action terminated the validity of the 1988-10-05 version and produced a new version, dated 2000-02-14 (with the new text), for the altered component.
@@ 3.4 Alteration Actions as Text Units
Although this alteration command already appears in the text of the caput (main section) of Art. 1º of Constitutional Amendment no. 26, as shown in Figure 14, generating a Text Unit for the Action entity structures and completes the information about the action/command, making explicit, for example, the validity periods of the text versions.
@@
The generated Text Unit is then associated with the Action entity in the graph. And it receives its own embedding, allowing, besides Version Text Units embeddings, for more sophisticated analyses of legislative evolution, facilitates the traceability of normative alterations, and improves the understanding of the life cycle of legal norms in the knowledge graph [14], in the context of RAG.
@@ 3.5 Metadata as Text Units
For example, for the Norm entity representing the legal norm “Lei nº 10.406 de 10/01/2002” (Law no. 10,406 of 2002-01-10), a Text Unit would be generated textually describing the entity’s properties [^22] This Text Unit was generated from the structured metadata registered in the knowledge graph of the ”Sistema de Gestão de Normas Jurídicas (Sigen)” (Legal Norm Management System) of the Federal Senate of Brazil. This metadata can also be viewed publicly, in structured form, through the Normas.leg.br portal, as in the example: https://normas.leg.br/?urn=urn:lex:br:federal:lei:2002-01-10;10406. Accessed on 2025-03-22.:
@@
This approach aligns with the concept of “Multi-vector embeddings[15], which proposes using multiple vectors to represent information at different levels of detail. Instead of generating a single, complete Text Unit for an entity encompassing all its properties and relationships, this work proposes generating one Text Unit for the entity’s intrinsic properties (general metadata), as exemplified above, and a separate Text Unit for each of its relationships.
@@
These metadata Text Units about the norm, or component, can be generated externally (from other information sources) and inserted directly into the graph, or be generated automatically from the properties and/or relationships of the entity in the graph [^24] As these properties (publication date, alternative names, level in the legal system, etc.) and relationships (correlation, succession, etc.) generally cannot be directly derived from the texts of the legal norms, they must be populated in the graph either manually or through semi-automatic extraction processes using other information sources. - although the original metadata might be structured in the graph, converting them to textual format allows their uniform integration and processing within the Graph RAG workflow.
@@
We can see these Text Units derived from metadata, additional to those derived from the texts of the norms/components, in Figure 16. The insertion of this type of Text Unit brings additional information to the Knowledge Graph and, consequently, to the vector database of contexts.
Figure 16: Knowledge graph illustrating the Text Units derived from both the texts of the norms and their metadata.
@@ 3.6 Hierarchical Grouping as Communities
The themes would then be inserted as Entities/Nodes Theme in the Knowledge Graph, each having an associated Text Unit with its description, which would be produced by a human domain expert, ensuring the curation of the produced texts [^26] An alternative would be to use an LLM to generate the theme descriptions, producing summaries from the descriptions of the sub-themes or, in the case of leaf themes, from the summaries of the associated norms. However, this approach has disadvantages: besides requiring intensive processing and high costs (considering that a theme can cover hundreds of norms), the summary’s effectiveness would heavily depend on the quality of the algorithms and models used..
@@ 3.7 Retrieval considering Structure
Another advantage of using a graph to represent structural entities, reflecting the hierarchical structure of legal texts, is that the user could choose the scope of their query (level of aggregation), for example, by defining that they wish to query about a specific theme (which would encompass all its sub-themes and associated norms, with their components and versions), or about a specific title of a norm (which would encompass all its subdivisions and versions), or about a specific article (which would encompass its subdivisions and versions), or about a specific version of a norm or a component (which would encompass its child versions), as presented in Figure 19.
@@
This choice would define a filter on the hierarchical index of the Text Units (and vectors/embeddings) to be retrieved, making the context retrieval for the response more focused. The hierarchical structure of the graph would then allow for more precise semantic navigation between different levels of detail. This “semantic navigation” manifests in the user’s ability to direct the search to a specific level of the legal hierarchy (theme, title, chapter, article, etc.), filtering the search space and focusing retrieval on the Text Units most relevant to their interest at that level of detail. Instead of a generic search across the entire graph, the user can explore legal knowledge in a more structured and hierarchy-guided manner.
@@ 4 Conclusion and Future Work
Aiming to enhance the analysis and comprehension of legal norms in the era of Artificial Intelligence, this article presented a detailed proposal for the construction of a legal knowledge graph, enriched with contextually relevant Text Units. The central objective was to significantly expand the scope of queries regarding legal norms, allowing Artificial Intelligence systems to address questions ranging from literal text retrieval to complex inquiries requiring an understanding of the hierarchical structure, the intrinsic semantics, the intra- and inter-relations and the temporal evolution of law.
Incorporating such structural and temporal knowledge is expected to significantly enhance the precision of legal AI responses, thereby reducing contextual errors and the risk of citing incorrect versions of the law.
@@
  • A detailed proposal for building a legal knowledge graph, integrating the predefined hierarchical structure, temporal versioning, and cumulative and contextually enriched Text Units, with an initial focus on structural entities.
@@
  • The proposal to generate specific Text Units for Alteration Actions as Text Units (e.g., repeal, text amendment), making the norm-modifying events themselves and their effects semantically retrievable.
  • The generation of Text Units from “structured Metadata” (intrinsic properties and relationships such as succession or correlation), allowing non-explicitly textual information in norms to be included in vector search and aligning with concepts like multi-vector embeddings.
@@
The main innovation lies in the adaptation and expansion of Graph RAG to the legal field, with a particular focus on structural entities, considered essential for identifying robust legal grounds. Semantic segmentation guided by the hierarchical structure, combined with the representation of cross-references, versioning and grouping into thematic communities, provides a natural and rich contextual basis for legal RAG strategies. The proposal to use the hierarchy of communities and themes designed by domain experts reflects the original intent of legislators and ensures informational curation, an essential aspect for the reliability of AI systems in law, in addition to minimizing the need for complex and costly graph construction processes.
@@
  • The proposed hierarchical structure for legal norms aligns with the Placing Graph approach of Bigraphs [18], which organizes entities into contextual hierarchies (e.g., temporal, thematic). Furthermore, the Linking Graph approach (hypergraph) could enrich the modeling by allowing multidimensional associations between components of norms, such as cross-references between articles or case law. As future work, it is proposed to explore the integration of Bigraphs to unify the hierarchical and associative modeling of legal norms. This approach could enhance the representation of spatio-temporal relationships (e.g., effectiveness of norms in different jurisdictions) and allow dynamic memory operations, such as automatic consolidation of versions or the identification of normative conflicts via semantic link analysis.
@@
Ultimately, by integrating structure, cross-references, temporality, and semantics into a legal knowledge graph, this work offers a significant step towards more intelligent, transparent, and effective Artificial Intelligence systems for the complex and fundamental domain of law. By facilitating precise information retrieval and making legislative analysis more accessible through a navigable and contextualized representation, this approach has the potential to democratize access to legal information and empower professionals and citizens in an increasingly regulated world.

### v2

@@ 3.2 Versions as Aggregations
In Figure 11, we observe that on the norm’s original date (1988-10-05), the original version of the component "tit2" is formed by the original versions of its children. That is, the text of "Título II” (Title II) on that date consists of it own text and the texts of its children from that same date.
On 1993-09-14, a new version for “tit2 cap4" was produced, either by its own alteration or indirectly by the alteration of one or more components in its tree. A more common implementation would be to produce new versions for the other children of "tit2" on 1993-09-14 and then produce a 1993-09-14 version for "tit2", which would be a composition [^13] In UML terms, Composition represents a strong whole-part relationship where the ”part” belongs exclusively to one ”whole” and shares its lifecycle; if the whole is deleted, the part is too. In our initial comparison, this would mean a parent version strictly owns child versions from the exact same date. of its children’s versions on the same date - similar to the formation of the "tit2" text on the original date. This would imply that, for each date on which the text of the norm was altered, a complete version of the norm would be produced from the composition of the complete versions of each immediate child of its articulation.
@@
We can verify the formation of these version aggregations in Figure 11, where new versions of the component "tit2" were produced as a consequence of new versions of one or more of its children on each date, reusing unchanged versions of its children where appropriate.
@@
When a version of the child component Cx does not exist for a date Tn, the most recent version of Cx, prior to Tn, will be used in the aggregation, as exemplified by V1 of C3, which is being used in the formation of version V2 of C1 (examples of this situation also appear in Figure 11). In this case, Version V1 of component C3 contributes to versions V1 and V2 of component C1, demonstrating that a child version can participate in multiple parent versions in different temporal contexts. The orange arrows indicate the aggregation relationship (Figure 12).
@@
In the context of RAG, it allows the system to retrieve not only the specific version of a component, but also the context of its parent and child versions, providing a more holistic and precise understanding of the norm at a given moment, facilitating more complex queries about the temporal evolution of norms [2].
@@ 3.3 Additional Context for Versions
For example, consider the caput of Art. 6º (presented in Figure 5) of the Brazilian Federal Constitution of 1988, signed and published on 1988-10-05. Thus, at the time of the Constitution’s publication, the Additional Context property of its original version (1988-10-05) will contain information about the start of its validity:
@@
For example, with the publication, on 2000-02-15, of “Emenda Constitucional nº 26 de 14/02/2000” (Brazilian Constitutional Amendment No. 26), signed on 2000-02-14, an action was commanded, through the caput of its Art. 1º, to alter the wording of the caput provision of Art. 6° of the Brazilian Federal Constitution of 1988 – the action is represented in the graph [^16] The modeling of the “Action presented in Figure 13 was simplified for didactic purposes. It can be improved using ontology techniques, such as those proposed in UFO (Unified Foundational Ontology), found at https://nemo.inf.ufes.br/projetos/ufo/. Accessed on 22.03.2025. by the Action entity (in brown in Figure 13). This action terminated the validity of the 1988-10-05 version and produced a new version, dated 2000-02-14 (with the new text), for the altered component.
For a more precise representation in the Graph, specialized entity types, derived from Action, can be created for each type of action between norms found in the treated legal system. For example, in the Brazilian case, types like "Alteração de Texto" (Text Change Action), "Renumeração" (Renumbering Action), "Revogação" (Repeal Action) etc. Thus, the entity representing the action could be included in the graph using the specialized type.
@@ 3.4 Alteration Actions as Text Units
Although this alteration command already appears in the text of the caput of Art. 1º of Constitutional Amendment no. 26 (amending component/norm), as shown in Figure 14, generating a Text Unit for the Action entity structures and completes the information about the action/command, making explicit, for example, the validity periods of the text versions.
@@
The generated Text Unit is then associated with the Action entity in the graph. And it receives its own embedding, allowing, besides Version Text Units embeddings, for more sophisticated analyses of legislative evolution, facilitates the traceability of normative alterations, and improves the understanding of the life cycle of legal norms in the knowledge graph in the context of RAG [14].
@@ 3.5 Metadata as Text Units
As example, for the Norm entity representing the legal norm “Lei nº 10.406 de 10/01/2002” (Law no. 10,406 of 2002-01-10), a Text Unit would be generated textually describing the entity’s properties [^18] This Text Unit was generated from the structured metadata registered in the knowledge graph of the ”Sistema de Gestão de Normas Jurídicas (Sigen)” (Legal Norm Management System) of the Federal Senate of Brazil. This metadata can also be viewed publicly, in structured form, through the Normas.leg.br portal, as in the example: https://normas.leg.br/?urn=urn:lex:br:federal:lei:2002-01-10;10406. Accessed on 2025-03-22.:
@@
This approach aligns with the concept of “Multi-aspect embeddings[19], which proposes using multiple vectors to represent information at different levels of detail or different aspects. Instead of generating a single, complete Text Unit for an entity encompassing all its properties and relationships, this work proposes generating one Text Unit for the entity’s intrinsic properties (general metadata), as exemplified above, and a separate Text Unit for each of its relationships.
@@
These metadata Text Units about the norm, or component, can be generated externally (from other information sources) and inserted directly into the graph, or be generated from the properties and/or relationships of the entity in the graph [^20] As these properties (publication date, alternative names, level in the legal system, etc.) and relationships (correlation, succession, etc.) generally cannot be directly derived from the texts of the legal norms, they must be populated in the graph either manually or through semi-automatic extraction processes using other information sources. - although the original metadata might be structured in the graph, converting them to textual format allows their uniform integration and processing within the Graph RAG workflow.
@@
We can see these Text Units derived from metadata, additional to those derived from versions of the norms/components, in Figure 16. The insertion of this type of Text Unit brings additional information to the Knowledge Graph and, consequently, to the vector database of contexts.
Figure 16: Knowledge graph illustrating the Text Units derived from both versions of the norms/components and their metadata.
@@ 3.6 Hierarchical Grouping as Communities
The themes would then be inserted as Entities/Nodes Theme in the Knowledge Graph, each having an associated Text Unit with its description, which would be produced by a human domain expert, ensuring the curation of the produced texts.
Another approach could be to use an LLM to generate the theme descriptions, producing summaries from the descriptions of the sub-themes or, in the case of leaf themes, from the summaries of the associated norms - it would require intensive processing and high costs (considering that a theme can cover hundreds of norms). The LLM would act as an assistant in a semi-automatic generation, where these descriptions would be reviewed and potentially updated by the human domain expert - the summary’s effectiveness would heavily depend on the quality of the algorithms and models used.
@@ 3.7 Retrieval considering Structure
Another advantage of using a graph to represent structural entities, reflecting the hierarchical structure of legal texts, is that the user could choose the scope of their query (level of aggregation), for example, by defining that they wish to query about a specific theme (which would encompass all its sub-themes and associated norms, with their components and versions), or about a specific title or article of a norm (which would encompass all its subdivisions and versions), or about a specific version of a norm or a component (which would encompass its child versions), as presented in Figure 19.
@@
The choice defines a filter on the hierarchical index of the Text Units (and vectors/embeddings) to be retrieved, making the context retrieval for the response more focused. The hierarchical structure of the graph allows for more precise semantic navigation between different levels of detail. This “semantic navigation” manifests in the user’s ability to direct the search to a specific level of the legal hierarchy (theme, title, chapter, article, etc.), filtering the search space and focusing retrieval on the Text Units most relevant to their interest at that level of detail. Instead of a generic search across the entire graph, the user can explore legal knowledge in a more structured and hierarchy-guided manner.
@@ 4 Conclusion and Future Work
Aiming to enhance the analysis and comprehension of legal norms in the era of Artificial Intelligence, this article presented a detailed proposal for the construction of a legal knowledge graph, enriched with contextually relevant Text Units. The central objective was to expand the scope of queries regarding legal norms, allowing Artificial Intelligence systems to address questions ranging from literal text retrieval to complex inquiries requiring an understanding of the hierarchical structure, the intrinsic semantics, the intra- and inter-relations and the temporal evolution of law.
Incorporating such structural and temporal knowledge is expected to enhance the precision of legal AI responses, thereby reducing contextual errors and the risk of citing incorrect versions of the law.
@@
  • A detailed proposal for building a legal knowledge graph, integrating the hierarchical structure, temporal versioning, and cumulative and contextually enriched Text Units, with an initial focus on structural entities.
@@
  • The proposal to generate specific Text Units for Alteration Actions (e.g., repeal, text amendment), making the norm-modifying events themselves and their effects semantically retrievable.
  • The generation of Text Units from structured Metadata (intrinsic properties and relationships such as succession or correlation), allowing non-explicitly textual information in norms to be included in vector search and aligning with concepts like multi-aspect embeddings.
@@
The main innovation lies in the adaptation and expansion of Graph RAG to the legal field, with a particular focus on structural entities, considered essential for identifying robust legal grounds. Semantic segmentation guided by the hierarchical structure, combined with the representation of cross-references, versioning and grouping into thematic communities, provides a natural and rich contextual basis for legal RAG strategies.
The proposal to use the hierarchy of communities and themes designed by domain experts reflects the original intent of legislators and ensures informational curation, an essential aspect for the reliability of AI systems in Law, in addition to minimizing the need for complex and costly graph construction processes.
@@
  • Integrating Bigraphs [18] with the proposed Graph RAG model could further enhance the representation of spatio-temporal relationships and enable dynamic memory operations, such as automatic version consolidation or the identification of normative conflicts via semantic link analysis.
@@
Ultimately, by integrating structure, cross-references, temporality, and semantics into a legal knowledge graph, this work offers a step towards more intelligent, transparent, and effective Artificial Intelligence systems for the complex and fundamental domain of Law. By facilitating precise information retrieval and making legislative analysis more accessible through a navigable and contextualized representation, this approach has the potential to democratize access to legal information and empower professionals and citizens in an increasingly regulated world.